Week of 200706

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Kate (chair, DB), Maarten (WLCG, ALICE), Ivan (ATLAS), Christoph (CMS), Darren (RAL), Dave M (FNAL), Andrew (TRIUMF), Xin (BNL), Pepe (PIC), Alberto A (monitoring), Gavin (computing), David B. (IN2P3)

Experiments round table:

  • ALICE -
    • Mostly business as usual.
    • NIKHEF VOBOX became unstable after update to latest SL6 kernel 2.6.32-754.6.3.el6
      • more and more VMEM allocated in the kernel, but not shown anywhere in ps
      • eventually the VM would become unresponsive and had to be rebooted
      • cured by downgrading to 2.6.32-696.30.1.el6
      • upgrade to CentOS 7 planned for Q3
      • we thank NIKHEF for debugging this over the weekend! (GGUS:147720)

Sites / Services round table:

  • ASGC: nc
  • BNL: NTR
  • CNAF: Scheduled Downtime: next Wednesday, July 8th, staring from 19.00 CEST, the general Internet connection to CNAF data center will experience several interruptions to allow the replacement of our frontier router with a new one.Network connectivity on LHC-OPN/ONE will not be directly affected: anyway, some external auxiliary services could be unable to reach the data center on the LHC-OPN/ONE network, thus causing problems also to data transfers on it or for remote data access from INFN T1 farm.The network connectivity should be completely restored at 24.00 CEST. GOCDB 29038
  • EGI: nc
  • FNAL: NTR
  • IN2P3: NTR
  • JINR: NTR
  • KISTI: nc
  • KIT: nc
  • NDGF: nc
  • NL-T1: nc
  • NRC-KI: nc
  • OSG: nc
  • PIC: For the upcoming Tuesday 14th of July, PIC will be in a complete scheduled downtime due to the following interventions: dCache upgrade, Enstore upgrade, HTCondor upgrade, Jupyter intervention in GPU platform. PIC will be in OUTAGE from 14-07-2020 08:00 (PIC local time) until 14-07-2020 14:00 (PIC local time). HTCondor will be stopped at the beginning of the SD. Additionally, we sent a message to the wlcg-data-archival list, since we have observed many positioning errors in our IBM LTO8 drives. We've been offered with new firmwares and things seem to improve. However, we would like to know about other site's experiences. Best regards,Manager on Duty
  • RAL: NTR
  • TRIUMF: NTR

AOB:

Edit | Attach | Watch | Print version | History: r19 < r18 < r17 < r16 < r15 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r19 - 2020-07-06 - GavinMcCance
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback